OpenStudy (tylermckinney16):

Suppose that under your health insurance policy, hospital expenses are subject to a $1,000 deductible and $250 per day copay. You get sick and are hospitalized for 4 days, and the bill (after insurance discounts are applied) comes to $9,000. How much of that hospital bill will you have to pay yourself?

OpenStudy (tylermckinney16):

@agent0smith

OpenStudy (agent0smith):

$1,000 deductible - you pay this $250 per day copay - multiply this by number of days Add those amounts together

OpenStudy (tylermckinney16):

$250*4=1000?

OpenStudy (agent0smith):

You also pay the deductible

OpenStudy (tylermckinney16):

So is it like this $250*4=1000 1000+1000=2000?

OpenStudy (agent0smith):

Yes

OpenStudy (tylermckinney16):

So thats the answer?